OpenStudy (anonymous):

kimes Consider the function f(x) = -5 x - 2 and find the following: b) The average rate of change between the points (a, f(a) ) and (b, f(b) ) . c) The average rate of change between the points (x, f(x) ) and (x+h, f(x+h) ) .

OpenStudy (jhannybean):

Already helped you find these.

OpenStudy (anonymous):

i know, can you check my answers?

OpenStudy (anonymous):

cause i got (-5b+5a)/b-a for b) and f for c)

OpenStudy (jhannybean):

\[b. \\ m=\frac{f(b)-f(a)}{b-a} = \frac{(-5b-2)-(-5a-2)}{b-a} = \frac{-5b+5a}{b-a}\qquad \checkmark \] \[c. \\m= \frac{f(x+h)-f(x)}{h} =\frac{-5(x+h)-2-(-5x-2)}{h} = \frac{-5x-5h-2+5x+2}{h} \\ m=\frac{-5h}{h} = -5 \qquad \color{red}{X} \]